The two plates of a capacitor hold +2500 μC and -2500 μC of charge, respectively, when the potential difference is 960V. What is the capacitance?


1
Expert's answer
2022-03-06T15:17:08-0500

The capacitance of the capacitor is given by

"C=\frac{q}{V}=\rm\frac{2500*10^{-6}\: C}{960\: V}=2.6*10^{-6}\: F=2.6\: \mu F"